# Conformal prediction of future insurance claims in the regression problem

### v2

@@ 2 Model-based approaches versus model-free approaches for prediction
$Y=g^{\star}(X)+\epsilon,\quad i=1,\ldots,n,$ (1)
@@
=$\mathcal{M}=\{(\mathcal{G},\mathcal{F})\},$
@@
where the lower bound $L(Z^{n})$ can be $\infty$- ∞ and the upper bound $U(Z^{n})$ can be +$\infty$+ ∞. If $\mathcal{G}$ and $\mathcal{F}$ both can be characterized by finitely many parameters, the model $\mathcal{M}$ is said to be a parametric model. For example, a linear model is a parametric model. If either $\mathcal{G}$ or $\mathcal{F}$ cannot be characterized by finitely many parameters, we say $\mathcal{M}$ is a non-parametric model. For example, if we take $\{g(x_{1},\ldots,x_{p})=a_{1}x_{1}+\ldots+a_{p}x_{p}\mid(a_{1},\ldots,a_{p})\in\mathbb{R}^{p}\}$ and =$\mathcal{F}=\{F\mid\text{$F$ is a distribution function supported on $\mathbb{R}$}\}$, then $\mathcal{M}$ will be a non-parametric model.
@@
The above two challenges in a model-based approach stem from the fact that prediction is conducted via a model. Therefore, one obvious way to circumvent these two difficulties is to perform prediction in one stroke: prediction is made from data directly without specifying a model. Such an approach is called a model-free approach for prediction. Technically, a model-free method for prediction is a mapping that maps data $Z^{n}$ to a prediction interval $(L(Z^{n}),U(Z^{n}))$. Figure 1 illustrates the difference between a model-based approach and a model-free approach. There are two paths from “Data” to “Prediction”. One first goes from “Data” to “Model” and then from “Model” to “Prediction”; this path denotes the model-based approach. The other path goes from “Data” directly to “Prediction”; it stands for the model-free approach.
@@
A distinguishing mark of a model-free approach is that the notion of a model never enters the scene. Many machine learning techniques, such as decision trees, K-nearest neighbors, and neural networks, share this feature. Some elementary estimators in non-parametric statistics, such as sample mean and sample variance, have the same feature too, though they are designed for estimation, not prediction. A model-free approach for prediction can be considered a prediction method in non-parametric statistics. However, as mentioned above, a non-parametric statistical model might still be subject to the adverse impact of the selection effect. Moreover, non-parametric predictive models (e.g., Jeon and Kim 2013, Hong and Martin 2017, and Richardson and Hartman 2018) can ensure asymptotically valid prediction at best, and none of them guarantees finite-sample validity. Finally, non-parametric prediction methods exist in the literature (e.g., Wilks 1941; Fligner and Wolfe 1976; Frey 2013, Hong and Martin 2021, Hong 2023). But these methods are all designed for unsupervised learning. Adopting them directly for prediction in the regression problem means we need to disregard the information provided by $X_{1},\ldots,X_{n}$, which is undesirable. Therefore, we need to seek a new method in supervised learning that is (i) model-free, (ii) tuning-parameter-free, and (iii) guarantees finite-sample validity.
@@ 3 Conformal prediction
1 Initialize: data $z^{n}=\{z_{1},\ldots,z_{n}\}$ and $x_{n+1}$, non-conformity measure $M$, and a grid of $y$ value;
2 for each possible $y$ do
@@ 5       Compute $\mathsf{pl}_{z^{n}}(x_{n+1},y)=(n+1)^{-1}\sum_{i=1}^{n+1}1\{\mu_{i}\geq\mu_{n+1}\}$;
6
7 end for
8Return $\mathsf{pl}_{z^{n}}(x_{n+1},y)$ for each $y$ on the grid;
@@
$1_{E}(x)=\left\{\begin{array}[]{ll}1,&\hbox{if $x\in E$,}\\ 0,&\hbox{otherwise.}\\ \end{array}\right.$
The quantity $\mu_{i}$, called the $i$-th non-conformity score, assigns a numerical value to $z_{i}$ to indicate how much $z_{i}$ agrees with the data in the $i$-th augmented bag $\widetilde{B}_{i}=z^{n}\cup\{z_{n+1}\}\backslash\{z_{i}\}$, where $z_{i}$ itself is excluded to avoid biases as in leave-one-out cross-validation. The function $\mathsf{pl}_{z^{n}}$ is called the plausibility function. It outputs a value between 00 and $1$, based on all non-conformity scores. The output of $\mathsf{pl}_{z^{n}}$ shows how plausible $z$ is a value of $Z_{n+1}$, based on the available data $Z^{n}=z^{n}$ and $X_{n+1}=x_{n+1}$. Using the output of the plausibility function $\mathsf{pl}_{z^{n}}$, the actuary can construct a $100(1-\alpha)\%$ conformal prediction region
@@
where $10<\alpha<$ 1. The next theorem establishes the finite-sample validity of the conformal prediction region in (2).
@@
If $\mathsf{P}$ denotes the distribution (i.e., the law) of an exchangeable sequence $Z_{1},Z_{2},\ldots$, then write $\mathsf{P}^{n+1}$ for the corresponding joint distribution of $Z^{n+1}=\{Z_{1},\ldots,Z_{n},Z_{n+1}\}$. For $10<\alpha<$ 1, define $t_{n}(\alpha)=(n+1)^{-1}\lfloor(n+1)\alpha\rfloor$, where $\lfloor a\rfloor$ denotes the greatest integer less than or equal to $a$. Then
@@
The proof is similar to its counterpart in the unsupervised learning context (e.g., Hong and Martin 2021). In great detail, $\mu_{1},\ldots,\mu_{n},\mu_{n+1}$ are exchangeable because $Z_{1},Z_{2},\ldots$ are exchangeable. Hence, the rank of $\mu_{n+1}$ follows the discrete uniform distribution on the set $\{1,\ldots,n,n+1\}$. By its definition, the plausibility function $\mathsf{pl}_{Z^{n}}(Z_{n+1})$ is proportional to the rank of $\mu_{n+1}$. Therefore, $\mathsf{pl}_{Z^{n}}(Z_{n+1})$ also follows the discrete uniform distribution on the set $\{1/(n+1),2/(n+1),\ldots,1\}$. For a given $10<\alpha<$ 1, if $(n+1)\alpha$ is an integer, then $\mathsf{P}^{n+1}\{\mathsf{pl}_{n+1}(Z_{n+1})\leq t_{n}(\alpha)\}=\alpha$. Otherwise, we will have $\mathsf{P}^{n+1}\{\mathsf{pl}_{n+1}(Z_{n+1})\leq t_{n}(\alpha)\}\leq t_{n}(\alpha)<\alpha$. Therefore, (3) holds. ∎
Theorem 1 implies that the prediction region in (2) is finite-sample (joint) valid in the sense that
@@ 4 Proposed method for insurance applications
It is clear from Algorithm 1 and Theorem 1 that conformal prediction is a model-free method. In particular, finite-sample validity, given by (4), holds regardless of the non-conformity measure an actuary chooses. However, to apply conformal prediction, an actuary must overcome several challenges. First, the shape of the conformal prediction region $C_{\alpha}(x,Z^{n})$ depends on the choice of the non-conformity measure. In particular, when the response variable $Y\in\mathbb{R}_{+}$, there is no guarantee that the conformal prediction region $C_{\alpha}(x,Z^{n})$ is an interval of the form $[0,b)$. The shape of $C_{\alpha}(x,Z^{n})$ can be a disjoint union of several intervals; see, for instance, Section 2.3 of Vovk et al. (2005) and Lei et al. (2013). If this happens, the resulting conformal region will not be practically useful. Second, perhaps more importantly, a close scrutiny of the argument in Section 3 reveals that the finite-sample validity of $C_{\alpha}(x,Z^{n})$, expressed by (4), hinges on the requirement that Algorithm 1 is implemented for all possible $y$ of $Y$. However, an actuary can only consider finitely many $y$ values in practice. This means $C_{\alpha}(x,Z^{n})$ cannot be determined exactly in general, compromising finite-sample validity—one of the key selling points of conformal prediction. Hong and Martin (2021) and Hong (2023) both take some special non-conformity measures and prove that corresponding $C_{\alpha}(x,Z^{n})$ is equivalent to another prediction interval which can be exactly determined (i.e., determined in finitely many steps). This is why the conformal prediction intervals in these two papers are finite-sample valid. In general, finite-sample validity is only possible but not automatically guaranteed when the actuary applies conformal prediction. Finally, even for an approximate implementation of Algorithm 1, i.e., we only consider a grid of $y$ values, the computation required could be prohibitively expensive.
To address these challenges together, we consider the following non-conformity measure. Let $B=\{z_{1},\ldots,z_{n}\}$ be a bag of observations of size $n$. For $i=1,\ldots,n$, let $z_{i}=(x_{i1},\ldots,x_{ip},y_{i})$ be the $i$-th observations in $B$. That is, $x_{ij}$ denotes the $i$-th observation of the the $j$-th feature. Suppose $z=(x_{1},\ldots,x_{p},y)$ is a provisional value of $Z_{n+1}$. We define $M$ to be
@@
In (5), $\sum_{i=1}^{n}(y_{i}-\sum_{j=1}^{p}x_{ij}/n)$ is the only term that depends on the bag $B$. Informally, it measures the aggregate difference between the response $Y$ and the sum of the features per data point. However, it is worth underlining that we are not proposing a model like $Y=X_{1}/n+\ldots+X_{p}/n+\varepsilon$, where $\varepsilon$ is some random error term. Conformal prediction is a model-free method; hence, there is no model to consider. $M$ in (5) is just a non-conformity measure we propose to address the above challenges. This non-conformity measure possesses several desirable properties: (i) the resulting prediction region is an interval of form $[0,b)$, (ii) there is a closed-form formula for the resulting conformal prediction interval, making implementation simple, (iii) determination of the resulting conformal prediction interval is exact, guaranteeing finite-sample validity. Properties (i)—(iii) immediately follow from the following theorem.
@@
Suppose the non-conformity measure is given by (5). For $1\leq i\leq n$, let $W_{i}=Y_{i}+\sum_{i=1}^{p}(X_{(n+1)j}-X_{ij})/n$. Then for any $10<\alpha<$ 1, the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n+1};Z^{n})$ for $Y_{n+1}$ equals $(0,W_{(k)})$, where $k=\min\{n,\lfloor(n+1)(1-\alpha)+1\rfloor\}$.
@@
which amounts to saying $\mu_{i}\geq\mu_{n+1}$ if and only if $Y_{n+1}\leq Y_{i}+(_{n+1}-S_{i})/n=W_{i}$ for all $1\leq i\leq n$. Therefore, $Y_{n+1}\leq W_{(k)}$ implies that $Y_{n+1}$ is less than at least $(n+1)-k+1$ $W_{i}$’s. Now the theorem follows from the definitions of $\mathsf{pl}_{z^{n}}(x,y)$ and $C_{\alpha}(x;Z^{n})$. ∎
@@ 1. $\bullet$
    ${\sf Bern}(p)$ denotes the Bernoulli distribution with success probability $10<p<$ 1;
@@ 3. $\bullet$
[…]
@@ 4. $\bullet$
$Y=X_{1}+\varepsilon,$ (7)
where $X_{1}\sim{\sf Pareto}(3,10)$, $\varepsilon\sim{\sf Gamma}(0.04,2.5)$, and $X_{1}$ and $\varepsilon$ are independent. All parameters are chosen to be the same as in Example 1: $\alpha=0.5\%$, $N=2,000$, $n=200$, $B=500$, and the number of trees in random forest is $500$. In this case, the oracle $99.5\%$ prediction interval is unknown, but we can approximate it as $[0,\hat{b})$, where $\hat{b}$ is the $99.5\%$ empirical quantile. Table 2 summarizes the simulation results.
@@
Similar to what we observed in Example 2, conformal prediction is the only method that achieves the nominal coverage probability level. However, this time conformal prediction intervals are more conservative. This is due to the fact that the data here has a higher degree of uncertainty than before. Note that $\mathsf{E}[X_{1}]$, $\mathsf{E}[X_{2}]$, and $\mathsf{E}[\epsilon]$ are $8$, $3.08$, and $0.5$, respectively. The second term in (8), being either $3$ or 00, adds a nontrivial layer of uncertainty to the data. Therefore, the conservativeness of the conformal prediction intervals is the price we must pay for the finite-sample validity. Indeed, all three methods behave more conservatively than in Example 2. The performance of random forest prediction intervals is similar to what we saw in the previous two examples: they tend to be aggressive in length, but, on average, they under-cover. The gamma regression model with the identity link is still a misspecified model. Model misspecification again wreaks havoc on prediction: the resulting GLM prediction intervals are too short to achieve the nominal coverage probability level.

### v3

@@ 2 Model-based approaches versus model-free approaches for prediction
$Y=g^{\star}(X)+\epsilon,$ (1)
@@
$\mathcal{M}=\{(\mathcal{G},\mathcal{F})\},$
@@
[…]
@@
The above two challenges in a model-based approach stem from the fact that prediction is conducted via a model. Therefore, one obvious way to circumvent these two difficulties is to perform prediction in one stroke: prediction is made from data directly without specifying a model. Such an approach is called a model-free approach for prediction. Technically, a model-free method for prediction is a mapping that maps data $Z^{n}$ directly to a prediction interval $(L(Z^{n}),U(Z^{n}))$ without training a model. Figure 1 illustrates the difference between a model-based approach and a model-free approach. There are two paths from “Data” to “Prediction”. One first goes from “Data” to “Model” and then from “Model” to “Prediction”; this path denotes the model-based approach. The other path goes from “Data” directly to “Prediction”; it stands for the model-free approach.
@@
A distinguishing mark of a model-free approach is that the notion of a model never enters the scene. Many machine learning techniques, such as decision trees and $K$-nearest neighbors (e.g., Trevor et al. 2009), share this feature. Some elementary estimators in non-parametric statistics, such as sample mean and sample variance, have the same feature too, though they are designed for estimation, not prediction. A model-free approach for prediction can be considered a prediction method in non-parametric statistics. However, as mentioned above, a non-parametric statistical model might still be subject to the adverse impact of the selection effect. Moreover, non-parametric predictive models (e.g., Jeon and Kim 2013, Hong and Martin 2017, and Richardson and Hartman 2018) can ensure asymptotically valid prediction at best, and none of them guarantees finite-sample validity. Finally, non-parametric prediction methods exist in the literature (e.g., Wilks 1941; Fligner and Wolfe 1976; Frey 2013, Hong and Martin 2021, Hong 2023). But these methods are all designed for unsupervised learning. Adopting them directly for prediction in the regression problem means we need to disregard the information provided by $X_{1},\ldots,X_{n}$, which is undesirable. Therefore, we need to seek a new method in supervised learning that is (i) model-free, (ii) tuning-parameter-free, and (iii) guarantees finite-sample validity.
We conclude this section with a few words on finite-sample validity and asymptotic validity of a prediction region. Suppose $\mathsf{P}$ is the distribution of $Z_{1}=(X_{1},Y_{1})$, and $D_{\alpha}(X_{n+1},Z^{n})$ is a $100(1-\alpha)\%$ prediction region (not necessarily constructed using conformal prediction) based on observations $Z^{n}$ and $X_{n+1}$. One of the key criteria for assessing the quality of $D_{\alpha}(X_{n+1},Z^{n})$ is its validity. We say $D_{\alpha}(X_{n+1},Z^{n})$ is asymptotically valid if
$\lim_{n\rightarrow\infty}\mathsf{P}^{n+1}\{Y_{n+1}\in D_{\alpha}(X_{n+1};Z^{n})\}\geq 1-\alpha\quad\text{for all $\mathsf{P}$},$ (2)
where $\mathsf{P}^{n+1}$ is the joint distribution for $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n}),(X_{n+1},Y_{n+1})$. Informally, asymptotic validity means if the actuary could sample from the true data-generating mechanism (1) forever, then eventually his coverage probability of his prediction region would be no less than the advertised confidence level $1-\alpha$, regardless of the distribution $\mathsf{P}$ of $(X_{1},Y_{1})$. However, the actuary can only have a finite sample in practice, and asymptotic validity does not tell us the sample size $n$ needed to ensure that the coverage probability of the prediction region $D_{\alpha}(X_{n+1},Z^{n})$ is no less than $1-\alpha$. For this reason, it would be practically more desirable if the coverage probability of $D_{\alpha}(X_{n+1},Z^{n})$ is at least $1-\alpha$ for any sample size $n$ and any distribution $\mathsf{P}$, that is,
$\mathsf{P}^{n+1}\{Y_{n+1}\in D_{\alpha}(X_{n+1};Z^{n})\}\geq 1-\alpha\quad\text{for all $(n,\mathsf{P})$},$ (3)
where $\mathsf{P}^{n+1}$ is the joint distribution for $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n}),(X_{n+1},Y_{n+1})$. If a prediction region $D_{\alpha}(X_{n+1},Z^{n})$ satisfies (3), we say it is finite-sample validity. In this case, given any sample of size $n$ and a confidence level $1-\alpha$, the actuary can construct a $100(1-\alpha)\%$ prediction region $D_{\alpha}(X_{n+1},Z^{n})$ satisfying (3), irrespective of what distribution $\mathsf{P}$ is.
@@ 3 Conformal prediction
1Initialize: data $z^{n}=\{z_{1},\ldots,z_{n}\}$ and $x_{n+1}$, non-conformity measure $M$, and a possible $y$ value;
@@ 4 Compute $\mathsf{pl}_{z^{n}}(x_{n+1},y)=(n+1)^{-1}\sum_{i=1}^{n+1}1\{\mu_{i}\geq\mu_{n+1}\}$;
5 Return $\mathsf{pl}_{z^{n}}(x_{n+1},y)$;
@@
[…]
@@
where $0<\alpha<1$. The next theorem establishes the finite-sample validity of the conformal prediction region in (4).
@@
If $\mathsf{P}$ denotes the distribution (i.e., the law) of an exchangeable sequence $Z_{1},Z_{2},\ldots$, then write $\mathsf{P}^{n+1}$ for the corresponding joint distribution of $Z^{n+1}=\{Z_{1},\ldots,Z_{n},Z_{n+1}\}$. For $0<\alpha<1$, define $t_{n}(\alpha)=(n+1)^{-1}\lfloor(n+1)\alpha\rfloor$, where $\lfloor a\rfloor$ denotes the greatest integer less than or equal to $a$. Then
@@
The proof is similar to its counterpart in the unsupervised learning context (e.g., Hong and Martin 2021). In great detail, $\mu_{1},\ldots,\mu_{n},\mu_{n+1}$ are exchangeable because $Z_{1},Z_{2},\ldots$ are exchangeable. Hence, the rank of $\mu_{n+1}$ follows the discrete uniform distribution on the set $\{1,\ldots,n,n+1\}$. By its definition, the plausibility function $\mathsf{pl}_{Z^{n}}(Z_{n+1})$ is proportional to the rank of $\mu_{n+1}$. Therefore, $\mathsf{pl}_{Z^{n}}(Z_{n+1})$ also follows the discrete uniform distribution on the set $\{1/(n+1),2/(n+1),\ldots,1\}$. For a given $0<\alpha<1$, if $(n+1)\alpha$ is an integer, then $\mathsf{P}^{n+1}\{\mathsf{pl}_{n+1}(Z_{n+1})\leq t_{n}(\alpha)\}=\alpha$. Otherwise, we will have $\mathsf{P}^{n+1}\{\mathsf{pl}_{n+1}(Z_{n+1})\leq t_{n}(\alpha)\}\leq t_{n}(\alpha)<\alpha$. Therefore, (5) holds. ∎
Theorem 1 implies that the prediction region in (4) is finite-sample (joint) valid:
@@ 4 Proposed method for insurance applications
It is clear from Algorithm 1 and Theorem 1 that conformal prediction is a model-free method. In particular, finite-sample validity, given by (6), holds regardless of the non-conformity measure an actuary chooses. However, to apply conformal prediction, an actuary must overcome several challenges. First, the shape of the conformal prediction region $C_{\alpha}(x,Z^{n})$ depends on the choice of the non-conformity measure. In particular, when the response variable $Y\in\mathbb{R}_{+}$, there is no guarantee that the conformal prediction region $C_{\alpha}(x,Z^{n})$ is an interval of the form $[0,b)$. The shape of $C_{\alpha}(x,Z^{n})$ can be a disjoint union of several intervals; see, for instance, Section 2.3 of Vovk et al. (2005) and Lei et al. (2013). If this happens, the resulting conformal region will not be practically useful. Second, perhaps more importantly, a close scrutiny of the argument in Section 3 reveals that determination of the $100\%$ conformal prediction region $C_{\alpha}(x;Z^{n})$ requires one to run Algorithm 1 for all possible $y$ value of $Y$. However, in practice an actuary can only run Algorithm 1 for finitely many $y$ values. This means $C_{\alpha}(x,Z^{n})$ cannot be determined exactly in general, compromising finite-sample validity—one of the key selling points of conformal prediction. Hong and Martin (2021) and Hong (2023) both take some special non-conformity measures and prove that corresponding $C_{\alpha}(x,Z^{n})$ is equivalent to another prediction interval which can be exactly determined (i.e., determined in finitely many steps). This is why the conformal prediction intervals in these two papers are finite-sample valid. In general, finite-sample validity is only possible but not automatically guaranteed when the actuary applies conformal prediction. Finally, even for an approximate determination of the conformal prediction region $C_{\alpha}(x;Z^{n})$, i.e., we only consider a grid of $y$ values, the computation required could be prohibitively expensive.
To address these challenges together, we consider the following non-conformity measure. Let $B=\{z_{1},\ldots,z_{n}\}$ be a bag of observations of size $n$. For $i=1,\ldots,n$, let $z_{i}=(x_{i1},\ldots,x_{ip},y_{i})$ be the $i$-th observations in $B$. That is, $x_{ij}$ denotes the $i$-th observation of the the $j$-th feature. Suppose $z=(x_{1},\ldots,x_{p},y)$ is a provisional value of a future observation to be predicted. We define $M$ to be
@@
In (7), $\sum_{i=1}^{n}(y_{i}-\sum_{j=1}^{p}x_{ij}/n)$ is the only term that depends on the bag $B$. Informally, it measures the aggregate difference between the response $Y$ and the sum of the features per data point. However, it is worth underlining that we are not proposing a model like $Y=X_{1}/n+\ldots+X_{p}/n+\varepsilon$, where $\varepsilon$ is some random error term. Conformal prediction is a model-free method; hence, there is no model to consider. $M$ in (7) is just a non-conformity measure we propose to address the above challenges. This non-conformity measure possesses several desirable properties: (i) the resulting prediction region is an interval of form $[0,b)$, (ii) there is a closed-form formula for the resulting conformal prediction interval, making implementation simple, (iii) determination of the resulting conformal prediction interval can be done in finitely many steps. Properties (i)—(ii) immediately follow from the following theorem. In addition, Theorem 2 shows the $100(1-\alpha)\%$ conformal conformal prediction region in this case is equivalent to an interval based on some order statistics, which can be determined in finitely many steps. Hence, there is no need to calculate the plausibility function $\mathsf{pl}_{Z^{n}}(X_{n+1},y)$ for all possible $y$.
@@
Suppose the non-conformity measure is given by (7). For $1\leq i\leq n$, let $W_{i}=Y_{i}+\sum_{i=1}^{p}(X_{(n+1)j}-X_{ij})/n$. Then for any $0<\alpha<1$, the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n+1};Z^{n})$ for $Y_{n+1}$ equals $(0,W_{(k)})$, where $k=\min\{n,\lfloor(n+1)(1-\alpha)+1\rfloor\}$.
@@
which amounts to saying $\mu_{i}\geq\mu_{n+1}$ if and only if $Y_{n+1}\leq Y_{i}+(S_{n+1}-S_{i})/n=W_{i}$ for all $1\leq i\leq n$. Therefore, $Y_{n+1}\leq W_{(k)}$ implies that $Y_{n+1}$ is less than at least (n+1)k+1(n+1)-k+1 $W_{i}$’s. Now the theorem follows from the definitions of $\mathsf{pl}_{z^{n}}(x,y)$ and $C_{\alpha}(x;Z^{n})$. ∎
@@ 1. $\bullet$
    ${\sf Bern}(p)$ denotes the Bernoulli distribution with success probability $0<p<1$;
@@ 3. $\bullet$
    where $\Gamma(\alpha)=\int_{0}^{\infty}t^{\alpha-1}e^{-t}dt$ is the Gamma function;
@@ 4. $\bullet$
$Y=X_{1}+\epsilon,$ (9)
[…]
@@
Similar to what we observed in Example 2, conformal prediction is the only method that achieves the nominal coverage probability level. However, this time conformal prediction intervals are more conservative. This is due to the fact that the data here has a higher degree of uncertainty than before. Note that $\mathsf{E}[X_{1}]$, $\mathsf{E}[X_{2}]$, and $\mathsf{E}[\epsilon]$ are $8$, $3.08$, and $0.5$, respectively. The second term in (10), being either $3$ or 0, adds a nontrivial layer of uncertainty to the data. Therefore, the conservativeness of the conformal prediction intervals is the price we must pay for the finite-sample validity. Indeed, all three methods behave more conservatively than in Example 2. The performance of random forest prediction intervals is similar to what we saw in the previous two examples: they tend to be aggressive in length, but, on average, they under-cover. The gamma regression model with the identity link is still a misspecified model. Model misspecification again wreaks havoc on prediction: the resulting GLM prediction intervals are too short to achieve the nominal coverage probability level.